OpenStudy (anonymous):

can someone solve this Suppose the population of a town is 8,200 and is growing 3% each year. Write an equation to model the population growth. Predict the population after 3 years.

OpenStudy (irishboy123):

startY1: population = 8200 end Y1: pop = P*(1+ 0.03) end Y2: pop = P*(1+ 0.03)*(1+ 0.03) see the pattern?